# Operationalizing Pluralistic Values in Large Language Model Alignment Reveals Trade-offs in Safety, Inclusivity, and Model Behavior

### v1

@@ Abstract
Although large language models (LLMs) are increasingly trained using human feedback for safety and alignment with human values, alignment decisions often overlook human social diversity. This study examines how incorporating pluralistic values affects LLM behavior by systematically evaluating demographic variation and design parameters in the alignment pipeline. We collect alignment data from US and German participants (N = 1,095 participants, 27,375 ratings) who rated LLM responses across five dimensions: Toxicity, Emotional Awareness (EA), Sensitivity, Stereotypical Bias, and Helpfulness. We fine-tuned multiple Large Language Models and Large Reasoning Models using preferences from different social groups while varying rating scales, disagreement handling methods, and optimization techniques. The results revealed systematic demographic effects: male participants rated responses 18% less toxic than female participants; conservative and Black participants rated responses 27.9% and 44% more emotionally aware than liberal and White participants, respectively. Models fine-tuned on group-specific preferences exhibited distinct behaviors. Technical design choices showed strong effects: the preservation of rater disagreement achieved roughly 53% greater toxicity reduction than majority voting, and 5-point scales yielded about 22% more reduction than binary formats; and Direct Preference Optimization (DPO) consistently outperformed Group Relative Policy Optimization (GRPO) in multi-value optimization. These findings represent a preliminary step in answering a critical question: How should alignment balance expert-driven and user-driven signals to ensure both safety and fair representation?
@@ Introduction
Prior studies have compared GPT-4’s safety annotations with human ratings across groups (Movva, Koh, and Pierson 2024), investigated cultural alignment via cross-lingual prompting and data mixtures (AlKhamissi et al. 2024), and introduced DICES (Aroyo et al. 2023), which frames disagreement as a signal for safety evaluation. These efforts lay the groundwork for understanding value diversity in alignment. Extending this work, we examine how demographic differences in feedback and design choices shape the collapse of pluralistic human values into a single model behavior during fine-tuning. Rather than implementing pluralistic alignment systems as proposed by Sorensen et al. (2024), we study how standard training homogenizes value variation and determines which group’s preferences dominate. Specifically, we ask: (1) How do models behave when aligned using feedback from different social groups? (2) How do technical choices such as rating scales, disagreement aggregation, and optimization methods affect learned values?
@@
To examine how social group differences and technical choices affect model alignment, we develop a bilingual alignment pipeline (English-German) with a focus on gender-related scenarios. Crucially, all participants, regardless of their own demographics, view the same set of gender-related prompt-response pairs, isolating the effects of rater demographics and design choices without confounding topic variation. The prompts are drawn from red-teaming, gender bias, and alignment (BeaverTails) benchmarks (Ganguli et al. 2022; Parrish et al. 2021; Ji et al. 2023), using gender-related keyword filters. Responses are generated using Wizard-Vicuna-7B-Uncensored-GPTQ (TheBloke 2023), selected because (1) its lack of safety filtering ensures exposure to a broad range of potentially problematic outputs, and (2) it is a stable open-source base model ensuring reproducibility. This yields 37,884 prompt-response pairs, from which we select 1,761 unique pairs for human evaluation. The pairs are translated into German using (DeepL 2024) and checked for semantic equivalence by a native speaker (see Supplementary A.1 and A.2).
@@
Participants rate model outputs along five alignment dimensions: Toxicity, Emotional Awareness, Sensitivity & Openness, Helpfulness, and Stereotypical Bias (see Supplementary A.3 for definitions provided to participants). Following recent sociotechnical alignment frameworks (Kirk et al. 2023), these dimensions capture how AI responses affect users, balancing ethical concerns (toxicity and stereotypical bias) with social dimensions (sensitivity, openness, and helpfulness) (Liu et al. 2024a; Bilquise, Ibrahim, and Shaalan 2022; Yin, Goh, and Hu 2024; Lissak et al. 2024; Ji et al. 2023).
This multidimensional approach expands beyond traditional helpfulness and harmfulness prompt–response pairs, yielding a total of 27,375 evaluations. Each dimension addresses aspects of human-AI interaction that different social groups may prioritize differently, making them relevant for studying value pluralism in alignment. Each dimension is rated using a 5-point Likert scale (from “Strongly Disagree” to “Strongly Agree”), allowing participants to express the intensity of their judgments rather than making binary classifications. Participants received clear definitions for each dimension (see supplementary A.3) before rating, ensuring consistent interpretation across cultural contexts.
@@
Experiment Setup
@@
The reference group is defined as White, US-based, Liberal, aged 18–30, and identifying as she/her/hers. Our analysis reveals several statistically significant demographic effects (see Table 1. Male participants rate responses as 18% less toxic and 20.9% less stereotypically biased compared to female participants. Conservative participants perceive responses as 27.9% more sensitive and 27% more emotionally aware than Liberals. Black or African American participants rate responses as 58% more sensitive and 44% more emotionally aware than White participants, a pattern consistent with work showing that Black Americans often attend more closely to emotional nuance and cultural sensitivity in AI interactions (Sandoval et al. 2025; Basoah et al. 2025). Additionally, participants aged 51-60 find responses 40.6% less helpful compared to younger participants (see Supplementary A.6).
@@
We conduct four fine-tuning experiments to examine how demographic variation and technical design choices affect model alignment. Experiments 1–3 rely on Direct Preference Optimization (DPO) (Rafailov et al. 2023), which learns from pairwise response comparisons. [^1] We use DPO for Experiments 1–3 because Experiment 4 shows it outperforms GRPO. Experiment 4 then contrasts DPO with Group-Relative Policy Optimization (GRPO) (Shao et al. 2024), an on-policy method that optimizes scalar rewards across multiple sampled completions using group-normalized advantages. Each method requires distinct data formatting, and we construct datasets accordingly. All experiments focus on toxicity and emotional awareness; these two dimensions are chosen to capture complementary safety concerns of harm avoidance and social understanding, where demographic and cultural backgrounds shape perceptions. Across the dataset, we obtain 5,475 ratings per dimension from 1,095 raters, applied to 1,761 unique prompt–response pairs, some of which receive multiple independent ratings (see Supplementary B.1).
@@
We fine-tune models on balanced subsets of human feedback to examine how demographic composition influences alignment outcomes. Three contrasts are tested: gender (female vs. male), political orientation (liberal vs. conservative), and ethnicity (White vs. Black).  [^2] Age affects helpfulness only; it shows no consistent effects for toxicity or emotional awareness, so we do not use it as a contrast. Each subgroup contributed an equal number of ratings to control for the size of the data set and distributional effects. Inter-annotator reliability, measured using Krippendorff’s $\alpha$ for toxicity annotations, was $\alpha{=}0.35$ (White) and $\alpha{=}0.44$ (Black) for ethnicity, $\alpha{=}0.39$ (Liberal) and $\alpha{=}0.33$ (Conservative) for political orientation, and $\alpha{=}0.38$ for both female and male raters, consistent with prior work (Ross et al. 2016; Bui, von der Wense, and Lauscher 2025); these values indicate substantial within-group variation rather than consensus. For each subgroup, a separate model is fine-tuned using DPO while preserving all five-point Likert ratings. The resulting models are evaluated on two alignment dimensions: Toxicity and Emotional Awareness, and compared against a baseline trained on feedback from conservative, male, and Black participants. Following Table 1, we focus on emotional awareness for the political spectrum and ethnicity, and on toxicity for gender, with complementary robustness checks (Supplementary B.6). This design enables systematic analysis of both within and cross-dimension effects. All experiments were replicated across seven model architectures to ensure robustness.
@@
To examine the effects of inter-annotator disagreement handling, we create five training datasets using different aggregation methods: (1) all data without aggregation, (2) majority vote, (3) full consensus (complete raters agreement only), (4) random selection (first rater’s rating), and (5) averaged ratings (mean of numeric-coded responses rounded to nearest category). We fine-tune separate DPO models on each dataset to assess how disagreement resolution strategies impact alignment performance. All models are trained using the 5-point Likert scale and evaluated on the Toxicity dimension.
@@
All experiments use LoRA fine-tuning Hu et al. (2022), with deterministic sampling (temperature=0.0) for reproducible evaluation across seven diverse model architectures (1B-14B parameters), ensuring consistent comparison of alignment effects across demographic groups and technical conditions. After fine-tuning the models on human feedback, we use GPT–4o–mini to score their toxicity and emotional-awareness outputs. We validate these LLM-generated scores against human expert judgements, achieving 85% agreement (see Supplementary B.2 and B.3). We use a DerSimonian–Laird random-effects meta-analysis (DerSimonian and Laird 1986) to pool effects across model architectures, accounting for both within-model variance and between-model heterogeneity. Full equations and derivations are provided in (Supplementary B.5).
@@ Results
Cross-Dimensional Rating Complexity
Participants frequently assign seemingly contradictory ratings to the same response across different alignment dimensions. By examining how often responses rated positively for one dimension were also rated positively for another, we found that 20.8% of emotionally aware responses were simultaneously rated as toxic, while 33.5% of emotionally aware responses were considered stereotypically biased. This pattern shows that participants can simultaneously perceive both positive and negative qualities in the same AI response, suggesting that alignment evaluation involves complex trade-offs rather than straightforward categorization (Figure 1). This aligns with evidence that users hold diverse, sometimes conflicting preferences rather than a single aggregated signal (Gölz, Haghtalab, and Yang 2025; Fleisig, Fazelpour et al. 2025).
Figure 1: Conditional probabilities between alignment dimensions, showing the likelihood of positive ratings in one dimension given positive ratings in another.
@@
Figure 2: Effects of demographic composition on fine-tuning outcomes. Significant model effects appear only in the left-hand comparisons. Purple circles show fine-tuned model effects: score differences between models trained on Liberal, Female, or White feedback versus Conservative, Male, or Black feedback. Blue triangles show survey-estimated demographic differences for the same groups. For Emotional Awareness (EA), positive values favour the first group; for Toxicity, negative values favour the first group. Error bars are 95% confidence intervals.
Demographic Feedback Shapes Specific Behaviors. Models fine-tuned on Liberal and White feedback produced higher emotional awareness scores than those trained on Conservative and Black feedback (pooled effects: 0.049, $p=0.010$, and 0.046, $p=0.001$, respectively). Similarly, Models fine-tuned on Female feedback also showed lower toxicity than those trained on Male feedback (pooled effect: 0.035-0.035, $p=0.002$). These effects were consistent across seven model architectures.
Effects Are Dimension-Specific. To test for generalization beyond the target dimensions, each demographically fine-tuned model was evaluated on both toxicity and emotional awareness. No statistically significant cross-dimensional effects were observed. For instance, the model fine-tuned on Female toxicity feedback did not significantly alter emotional awareness ($p=0.860$), while Liberal and White emotional-awareness models showed no effect on toxicity ($p=0.500$ and $p=0.880$, respectively).
These findings demonstrate that demographic composition produces measurable yet dimension-specific effects on alignment: subgroup value preferences are reliably encoded in fine-tuned models without introducing unintended behavioral shifts across unrelated alignment dimensions.
@@
We examine how the granularity of the rating scale affects alignment effectiveness by comparing 5-point, 3-point, and binary scales using data from the toxicity dimension and DPO fine-tuning. We examine scale granularity using toxicity ratings, concentrating on this dimension to keep the comparison clean and analytically focused (Figure 3).
Granular Scales Achieve Superior Alignment Performance
@@
Feedback Granularity Impacts Alignment Effectiveness
5-point scales are approximately 22% more effective than binary scales at reducing toxicity (effect sizes: 0.242-0.242 vs. 0.198-0.198); This pattern is consistent with findings that multi-level cardinal feedback (such as 5-point scale) yields more learnable reward models than pairwise preferences, while maintaining similar levels of inter-rater reliability (Kreutzer, Uyheng, and Riezler 2018). This performance gap reveals that reducing human feedback to binary choices for individual alignment tasks, systematically under-utilizes available preference information. While binary formats are widely adopted in preference learning, our findings demonstrate measurable costs to the effectiveness of alignment.
Figure 3: Rating Scale Effects on Toxicity Reduction. (A) All scales reduce toxicity vs. control (no fine-tuning), with the 5-point most effective. (B) 5-point scales significantly outperform binary. Error bars: 95% CIs. Lower values indicate reduced toxicity.
@@
We examine how different approaches to handling inter-annotator disagreement affect alignment outcomes by comparing five aggregation strategies: preserving all ratings, averaging ratings, majority vote, random selection, and requiring full consensus (Figure 4).
@@
Figure 4: Disagreement Handling Strategy Effects on Alignment Training. (A) Strategy performance vs. control. (B) Pairwise strategy comparisons. Error bars show 95% CIs. Lower values indicate better performance.
@@
Figure 5: DPO and GRPO Optimization Methods Comparison. (A-B) Performance of each DPO and GRPO. (C-D) Pairwise comparisons between single-objective and multi-objective DPO approaches. Error bars show 95% CIs.
@@
Safety judgments are not universal but reflect specific demographic perspectives. Male and female participants rated identical responses with an 18% difference in perceived toxicity, while Conservative and Black participants reported 27.9% and 44% higher emotional awareness ratings, respectively, compared to Liberal, White participants. Current alignment approaches that aggregate these differences away may systematically exclude safety-relevant perspectives. Preserving disagreement achieved the strongest toxicity reduction, outperforming the majority vote by 53%, and other strategies. This suggests alignment annotators’ “noise” may be an essential safety signal. Technical choices such as rating scales (5-point exceeding binary by 22%); and optimization methods (DPO exceeding GRPO) profoundly impact safety performance. Rather than trading off safety against inclusivity, we find that inclusive approaches enhance safety outcomes.
@@
It is important to note several limitations ( Supplementary A.5). Our dataset covers only two countries and has uneven demographic representation, with conservatives, gender minorities, and older adults underrepresented. Model evaluation used GPT-4o-mini, although its judgments were cross-validated against human reviewers with high agreement. Our optimization analysis focused on DPO and GRPO, leaving other methods for future work. These constraints do not affect the core patterns but highlight the need for more diverse datasets and broader evaluation of optimization approaches.

### v2

@@ Abstract
Although large language models (LLMs) are increasingly trained using human feedback for safety and alignment with human values, alignment decisions often overlook human social diversity. This study examines how incorporating pluralistic values affects LLM behavior by systematically evaluating demographic variation and design parameters in the alignment pipeline. We collect alignment data from US and German participants (N = 1,095 participants, 27,375 ratings) who rated LLM responses across five dimensions: Toxicity, Emotional Awareness (EA), Sensitivity, Stereotypical Bias, and Helpfulness. We fine-tuned multiple Large Language Models and Large Reasoning Models using preferences from different social groups while varying rating scales, disagreement handling methods, and optimization techniques. The results revealed systematic demographic effects: male participants rated responses 18% less toxic than female participants; conservative and Black participants rated responses 27.9% and 44% higher on EA than liberal and White participants, respectively. Models fine-tuned on group-specific preferences exhibited distinct behaviors. Technical design choices showed strong effects: the preservation of rater disagreement achieved roughly 53% greater toxicity reduction than majority voting, and 5-point scales yielded about 22% more reduction than binary formats; and, Direct Preference Optimization (DPO) consistently outperformed Group Relative Policy Optimization (GRPO) in multi-value optimization. These findings represent a preliminary step in answering a critical question: How should alignment balance expert-driven and user-driven signals to ensure both safety and fair representation?
@@ Introduction
Prior studies have compared GPT-4’s safety annotations with human ratings across groups (Movva, Koh, and Pierson 2024), investigated cultural alignment via cross-lingual prompting and data mixtures (AlKhamissi et al. 2024), and introduced DICES (Aroyo et al. 2023), which frames disagreement as a signal for safety evaluation. Extending this work, we examine how demographic differences in feedback and design choices shape the collapse of pluralistic human values into a single model behavior during fine-tuning. Rather than implementing pluralistic alignment systems as proposed by Sorensen et al. (2024), we study how standard training homogenizes value variation and determines which group’s preferences dominate. Specifically, we ask: (1) How do models behave when aligned using feedback from different social groups? (2) How do technical choices such as rating scales, disagreement aggregation, and optimization methods affect learned values?
@@
To examine how social group differences and technical choices affect model alignment, we develop a bilingual alignment pipeline (English-German) with a focus on gender-related scenarios. Crucially, all participants, regardless of their own demographics, view the same set of gender-related prompt-response pairs, isolating the effects of rater demographics and design choices without confounding topic variation. The prompts are drawn from red-teaming, gender bias, and alignment (BeaverTails) benchmarks (Ganguli et al. 2022; Parrish et al. 2021; Ji et al. 2023), using gender-related keyword filters. Responses are generated using Wizard-Vicuna-7B-Uncensored-GPTQ (TheBloke 2023), selected because (1) its lack of safety filtering ensures exposure to a broad range of potentially problematic outputs, and (2) it is a stable open-source base model ensuring reproducibility. This yields 37,884 prompt-response pairs, from which we select 1,761 unique pairs for human evaluation. The pairs are translated into German using DeepL (2024) and checked for semantic equivalence by a native speaker (see Supplementary A.1 and A.2).
@@
Participants rate model outputs along five alignment dimensions: Toxicity, Emotional Awareness, Sensitivity & Openness, Helpfulness, and Stereotypical Bias (see Supplementary A.3 for definitions provided to participants). Following recent sociotechnical alignment frameworks (Kirk et al. 2023), these dimensions capture how AI responses affect users, balancing ethical concerns (toxicity and stereotypical bias) with social dimensions (sensitivity, openness, and helpfulness) (Liu et al. 2024a; Bilquise, Ibrahim, and Shaalan 2022; Yin, Goh, and Hu 2024; Lissak et al. 2024; Ji et al. 2023). This multidimensional approach expands beyond traditional helpfulness and harmfulness prompt–response pairs, yielding a total of 27,375 evaluations. Each dimension addresses aspects of human-AI interaction that different social groups may prioritize differently, making them relevant for studying value pluralism in alignment. Each dimension is rated using a 5-point Likert scale (from “Strongly Disagree” to “Strongly Agree”), allowing participants to express the intensity of their judgments rather than making binary classifications. Participants received clear definitions for each dimension (see supplementary A.3) before rating, ensuring consistent interpretation across cultural contexts.
@@
Data Analyses Informing Experiment
@@
Cross-Dimensional Rating Complexity
Participants frequently assign seemingly contradictory ratings to the same response across different alignment dimensions. By examining how often responses rated positively for one dimension were also rated positively for another, we found that among responses rated as emotionally aware, 20.8% were also rated as toxic, and 33.5% were also rated as stereotypically biased. This pattern shows that participants can simultaneously perceive both positive and negative qualities in the same AI response, suggesting that alignment evaluation involves complex trade-offs rather than straightforward categorization (Figure 1). This aligns with evidence that users hold diverse, sometimes conflicting preferences rather than a single aggregated signal (Gölz, Haghtalab, and Yang 2025; Fleisig, Fazelpour et al. 2025).
Figure 1: Conditional probabilities between alignment dimensions. Each cell shows the probability of a positive rating in the column given a positive rating in the row.
@@
The reference group is defined as White, US-based, Liberal, aged 18–30, and identifying as she/her/hers. Our analysis reveals several statistically significant demographic effects (see Table 1). Male participants rate responses as 18% less toxic and 20.9% less stereotypically biased compared to female participants. Conservative participants perceive responses as 27.9% more sensitive and 27% more emotionally aware than Liberals. Black or African American participants rate responses as 58% more sensitive and 44% more emotionally aware than White participants, a pattern consistent with work showing that Black Americans often attend more closely to emotional nuance and cultural sensitivity in AI interactions (Sandoval et al. 2025; Basoah et al. 2025). Additionally, participants aged 51-60 find responses 40.6% less helpful compared to younger participants (see Supplementary A.6).
@@
We conduct four fine-tuning experiments to examine how demographic variation and technical design choices affect model alignment. Experiments 1–3 rely on Direct Preference Optimization (DPO) (Rafailov et al. 2023), which learns from pairwise response comparisons. We use DPO for Experiments 1–3 because Experiment 4 shows it outperforms GRPO. Experiment 4 then contrasts DPO with Group-Relative Policy Optimization (GRPO) (Shao et al. 2024), an on-policy method that optimizes scalar rewards across multiple sampled completions using group-normalized advantages. Each method requires distinct data formatting, and we construct datasets accordingly. All experiments focus on toxicity and emotional awareness; these two dimensions are chosen to capture complementary safety concerns of harm avoidance and social understanding, where demographic and cultural backgrounds shape perceptions. Across the dataset, we obtain 5,475 ratings per dimension from 1,095 raters, applied to 1,761 unique prompt–response pairs, some of which receive multiple independent ratings (see Supplementary B.1, B.2 and B.3). All experiments are replicated across seven model architectures to ensure robustness (see Supplementary Table 9 for model overviews).
@@
We fine-tune models on balanced subsets of human feedback to examine how demographic composition influences alignment outcomes. Three contrasts are tested: gender (female vs. male), political orientation (liberal vs. conservative), and ethnicity (White vs. Black). [^1] Age affects helpfulness only; it shows no consistent effects for toxicity or emotional awareness, so we do not use it as a contrast. Each subgroup contributed an equal number of ratings to control for the size of the data set and distributional effects. Inter-annotator reliability, measured using Krippendorff’s $\alpha$ for toxicity annotations, was $\alpha{=}0.35$ (White) and $\alpha{=}0.44$ (Black) for ethnicity, $\alpha{=}0.39$ (Liberal) and $\alpha{=}0.33$ (Conservative) for political orientation, and $\alpha{=}0.38$ for both female and male raters, consistent with prior work (Ross et al. 2016; Bui, von der Wense, and Lauscher 2025); these values indicate substantial within-group variation rather than consensus. For each subgroup, a separate model is fine-tuned using DPO while preserving all five-point Likert ratings. The resulting models are evaluated on two alignment dimensions: Toxicity and Emotional Awareness, and compared against a baseline trained on feedback from conservative, male, and Black participants. Following Table 1, we focus on emotional awareness for the political spectrum and ethnicity, and on toxicity for gender, with complementary robustness checks (Supplementary B.6). This design enables systematic analysis of both within and cross-dimension effects.
@@
To examine the effects of inter-annotator disagreement handling, we create five training datasets using different aggregation methods: (1) all data without aggregation, (2) majority vote, (3) full consensus (complete rater agreement only), (4) random selection (first rater’s rating), and (5) averaged ratings (mean of numeric-coded responses rounded to nearest category). We fine-tune separate DPO models on each dataset to assess how disagreement resolution strategies impact alignment performance. All models are trained using the 5-point Likert scale and evaluated on the Toxicity dimension.
@@
All experiments use LoRA fine-tuning (Hu et al. 2022), with deterministic sampling (temperature=0.0) for reproducible evaluation across seven diverse model architectures (1B-14B parameters), ensuring consistent comparison of alignment effects across demographic groups and technical conditions. After fine-tuning the models on human feedback, we use GPT–4o–mini to score their toxicity and emotional-awareness outputs. We validate these LLM-generated scores against human expert judgements, achieving 85% agreement (see Supplementary B.2 and B.3). We use a DerSimonian–Laird random-effects meta-analysis (DerSimonian and Laird 1986) to pool effects across model architectures, accounting for both within-model variance and between-model heterogeneity. Full equations and derivations are provided in (Supplementary B.5) B.
@@ Results
Figure 2: Effects of demographic composition on fine-tuning outcomes. Purple circles show model differences between groups (Liberal, Female, White minus Conservative, Male, Black, respectively) on Emotional Awareness or Toxicity ratings; blue triangles show survey-estimated differences for the same contrasts. Positive values indicate higher Emotional Awareness for the first group; negative values indicate lower Toxicity for the first group. Error bars show 95% confidence intervals.
Demographic Feedback Shapes Specific Behaviors.
Models fine-tuned on Liberal and White feedback produced higher emotional awareness scores than those trained on Conservative and Black feedback (pooled effects: 0.049, $p=0.010$, and 0.046, $p=0.001$, respectively). Similarly, models fine-tuned on Female feedback showed lower toxicity than those trained on Male feedback (pooled effect: 0.035-0.035, $p=0.002$). These effects were consistent across seven model architectures.
Effects Are Dimension-Specific.
To test for generalization beyond the target dimensions, each demographically fine-tuned model was evaluated on both toxicity and emotional awareness. No statistically significant cross-dimensional effects were observed. For instance, the model fine-tuned on Female toxicity feedback did not significantly alter emotional awareness ($p=0.860$), while Liberal and White emotional-awareness models showed no effect on toxicity ($p=0.500$ and $p=0.880$, respectively). These findings demonstrate that demographic composition produces measurable yet dimension-specific effects on alignment: subgroup value preferences are reliably encoded in fine-tuned models without introducing unintended behavioral shifts across unrelated alignment dimensions.
@@
We examine how the granularity of the rating scale affects alignment effectiveness by comparing 5-point, 3-point, and binary scales using data from the toxicity dimension and DPO fine-tuning. We examine scale granularity with a focus on toxicity ratings (Figure 3).
Granular Scales Improve Alignment Performance
@@
Implications of Scale Granularity for Model Learning
5-point scales are 22% more effective than binary scales at reducing toxicity (effect sizes: 0.242-0.242 vs. 0.198-0.198); this pattern is consistent with findings that multi-level cardinal feedback (such as 5-point scale) yields more learnable reward models than pairwise preferences, while maintaining similar levels of inter-rater reliability (Kreutzer, Uyheng, and Riezler 2018). This performance gap reveals that reducing human feedback to binary choices for individual alignment tasks systematically under-utilizes available preference information. While binary formats are widely adopted in preference learning, our findings show measurable costs to the effectiveness of alignment.
Figure 3: Rating Scale Effects on Toxicity Reduction. (A) All scales reduce toxicity relative to the control (no fine-tuning), with the 5-point most effective. (B) 5-point scales significantly outperform binary. Error bars: 95% CIs. Lower values indicate reduced toxicity.
@@
We examine how approaches to handling inter-annotator disagreement affect alignment outcomes by comparing five aggregation strategies: preserving all ratings, averaging, majority vote, random selection, and full consensus (Figure 4).
@@
Figure 4: Disagreement Handling Strategy Effects on Alignment Training. (A) Strategy performance relative to the control (no fine-tuning) on toxicity. (B) Pairwise strategy comparisons. Error bars show 95% CIs. Lower values indicate better performance.
@@
Figure 5: DPO and GRPO Optimization Methods Comparison. (A-B) Performance of each DPO and GRPO trained model. (C-D) Pairwise comparisons between single-objective and multi-objective DPO approaches. Error bars show 95% CIs.
@@
Safety judgments are not universal but reflect specific demographic perspectives. Male and female participants rated identical responses with an 18% difference in perceived toxicity, while Conservative and Black participants reported 27.9% and 44% higher emotional awareness ratings, respectively, compared to Liberal, White participants. Current alignment approaches that aggregate these differences away may systematically exclude safety-relevant perspectives. Preserving disagreement achieved the strongest toxicity reduction, outperforming the majority vote by 53%, and other strategies. This suggests alignment annotators’ “noise” may be an essential safety signal. Technical choices such as rating scales (5-point exceeding binary by 22%) and optimization methods (DPO exceeding GRPO) profoundly impact safety performance. Rather than trading off safety against inclusivity, we find that inclusive approaches enhance safety outcomes.
@@
We conclude with limitations (see more in Supplementary A.5). Our dataset covers only two WEIRD-dominant countries and has uneven demographic representation (underrepresenting conservatives, gender minorities, and older adults). Model evaluation used GPT-4o-mini, though its judgments are in high agreement with human reviewers. Our optimization analysis focused on DPO and GRPO, leaving other methods for future work. These constraints do not affect the core patterns but highlight the need for more diverse datasets and broader evaluation of optimization approaches.